### v1

@@ 3.2.3 Highlighting ROI
Segmenting the lesion regions from the skin cancer images and leveraging the segmented portion for classification typically results in a better performance [33, 34]. However, incorporating a segmentation model like U-Net remarkably increases the computational burden. On the other hand, relying on traditional segmentation techniques might produce inconsistent results. Therefore, the lesion regions in the skin cancer images are highlighted to ensure feature preservation and prioritize relevant information. This step encompasses three key phases: grayscale conversion, image segmentation, and linear blending.
@@
Following the grayscale conversion, the image is then segmented with Otsu’s method. Otsu’s method calculates a threshold value based on the pixel intensity of the image which is leveraged for binary segmentation [35]. For measuring the optimal threshold value, it iterates over the potential threshold values, and the value having the maximizing inter-class variance is returned as the threshold. The formula for inter-class variance is calculated using Equation 4. In this equation, $\omega_{0}(t)$ and $\omega_{1}(t)$ represent the probabilities of two classes separated by a threshold $t$. The equations of $\omega_{0}(t)$ and $\omega_{1}(t)$ are presented in Equation 5 and 6 respectively. $\omega_{0}(t)$ is the cumulative probabilities of pixel intensities up to the threshold $t$ while $\omega_{1}(t)$ measures the probabilities from $t+1$ to $L-1$, where $L-1$ is the highest intensity level. In Equation 4, $\sigma_{0}$ and $\sigma_{1}$ are the variances of two classes.
@@ 3.2.4 Unsharp Masking
Unsharp masking is an image sharpening method that employs low pass filters to sharpen details and improve overall visibility [36]. It improves skin cancer images by accentuating high-frequency components corresponding to edges and tiny details. The image sharpening algorithm has three major steps. In the first step, the input image is convoluted with a low pass filter. Let $I$ define the input image and $f$ be a low pass filter. Now the resulting blurred image, $I_{b}$ is obtained by operation illustrated in Equation 9. The selection of the low pass filter plays a crucial role in this step. For the proposed solution, the kernel is automatically determined from the standard deviation values of 2.0 in both the X and Y directions.
@@ 3.2.6 Resizing
Image resizing is also another frequently employed image preprocessing technique, particularly in the context of transfer learning. Typically state-of-the-art image classifiers are trained on ImageNet, that have a resolution of 469$\times$387 [37]. However, they are typically resized to 255$\times$255 or 224$\times$244 [38]. Therefore, resizing the images to 224$\times$224 typically yields a better performance. Moreover, lowering the image size significantly reduces the computation. Since the images in the Kaggle dataset are already resized to the specified size, only the images from the HAM1000 dataset undergo this stage.
@@ 3.4 Teacher Model Creation
We have constructed a robust teacher model made of three feature extractors for guiding the lightweight student model. As presented in Figure 4, the proposed teacher model is composed of two streams. The first stream is composed of two CNN feature extractors: ResNet152V2 [39] and ConvNeXtBase [40], two state-of-the-art image classifiers. Since the feature extraction process of transformer architecture is different from CNNs, the second stream incorporates ViT Base [9], a robust transformer network. The features from these streams are fused to construct a strong teacher model.
ResNet152V2, a variant of the ResNet architecture, is specifically engineered to mitigate the challenge of the vanishing gradient problem. This issue arises during the training of deep networks, where gradients diminish significantly during backpropagation, leading to slowed or stalled learning progress. Residual networks address this issue through the integration of residual blocks. A residual block creates a shortcut or residual connection by adding the block’s input to its output. The architectural enhancement allows the network to learn the residual mapping, which facilitates the model’s ability to detect gradients and identity mappings during training. Due to the fine gradient preservation, it is possible to create deep neural networks without encountering the vanishing gradient problem. ResNet152V2 extends the original ResNet architecture by incorporating batch normalization that speeds up the training process [41]. This model consists of 152 layers that effectively extract useful features from the image, making the model a reliable classifier.
@@
The first stream of the proposed teacher model concatenates the features extracted from the pretrained ResNet152V2 and ConvNeXtBase. The combined features are then passed to batch normalization and Squeeze and Excitation (SE) block. SE block is an attention mechanism that prioritizes the channels that are more important for classification [42]. The block consists of three operations: squeezing, exciting, and recalibrating. In the squeeze operation, illustrated in Equation 13, the spatial dimension of a channel is reduced with global average pooling. Let $X_{c}$ denote a feature map with $H$ height and $W$ width. Now, the summary of the channel, denoted with scaler $Z_{c}$, is obtained by taking the average of the channel $X_{c}$.
@@
The second stream, on the other hand, consists of a ViT Base and a batch normalization layer. The features extracted from the first stream and the second stream are combined to construct the final fusion model. The fusion model then follows three blocks of dropout and dense layer. The first two dropout layer randomly drops 50% neurons and the last layer drops 20% neurons. The dropout layers prevent overfitting. The first three dense layers, on the other hand, consist of 256, 128, and 64 neurons which are activated with Scaled Exponential Linear Unit (SELU). The SELU activation function, known for its simplicity and effectiveness in preventing overfitting, is utilized in the final layer, comprising two neurons employing the softmax activation function [43]. The final layer comprises two neurons utilizing the softmax activation function. The model has over 236 million parameters out of which 3.39 million parameters are trainable. The size of the model is 975.85 megabytes.
@@ 3.5 Student Model Creation
The model is trained with knowledge distillation, a technique in which a simple model, known as the student, is trained to replicate the predictions of a more complex model, known as the teacher model [14]. Figure 6 presents an overview of the knowledge distillation process. The approach involves sharing knowledge from the teacher to the student through softened probabilities, which are adapted using two key hyperparameters: temperature and alpha. The temperature parameter regulates the level of softening, which influences the smoothness of the probability distribution, whilst the alpha parameter sets the weighting of hard and soft targets throughout the training phase. This enables the student model to benefit from the rich knowledge stored in the teacher’s predictions, resulting in a more efficient and accurate learning process.
@@ 3.6 Ablation Study
In order to determine the optimal hyperparameters, an extensive ablation study has been conducted on the student model. The outcomes of the study on the HAM10000 and the Kaggle datasets are presented in Table 3 and 4 respectively. The results depict a similar impact of the seven hyperparameters. According to the analysis, the feature conversion mechanism is the most influential one among the seven parameters compared. While converting 2D future maps into 1D feature vectors, global max pooling produces the best performance in both the dataset while global average pooling has resulted in a poor performance. However, optimizer, temperature, alpha, and number of epochs are some of the least important hyperparameters. On the HAM10000 dataset, Adafactor has exhibited superior performance, outperforming RSMprop, which achieved the least accuracy. However, on the Kaggle dataset, RSMprop has emerged as the best optimizer among the four optimizers compared. Due to the conflicting outcome, Adafactor has been selected as the optimizer due to its faster convergence [44]. Another conflicting outcome is produced with the temperature hyperparameter. On the HAM10000 dataset, the best classification performance is produced with temperature 1, and temperature 5 results in the worst performance. On the other hand, the best performance is achieved with a temperature value of 5 on the Kaggle dataset. In the proposed model, the temperature is set to 1, which implies that the soft targets from the teacher model are passed without any scaling. On both datasets, an alpha value of 0.5, a learning rate of 0.002, and a batch size of 64 have produced the best performance. Notably, the learning rate has been one of the most influential hyperparameters in this experiment. In both datasets, a learning rate of 0.1 has created an overfitted model while reducing the learning rate has gradually improved the accuracy.
@@ 3.7 Model Quantization
Model quantization is a method that represents parameters and activations with a lower bit precision than the conventional 64-bit floating-point format, thereby reducing the memory footprint and computing demands of deep neural networks [45]. This procedure is essential for deploying models on devices with limited memory and processing capability, like smartphones or edge devices. While quantization increases efficiency, over-quantization can reduce model accuracy due to the loss of fine-grained data [46].
@@ 4.2 Evaluation Metrics
To evaluate the model’s performance, we have considered five evaluation metrics including accuracy, precision, recall, F1-score and Matthews Correlation Coefficient (MCC). While accuracy presents the overall correctness of the model, precision measures the accuracy of positive predictions. Recall, on the other hand, the ratio of true positive predictions to the actual positives. Accuracy measures the harmonic mean of precision and recall, presenting a more balanced result. MCC considers true positive, true negative, false positive, and false negative values to asses the models’s performance. Since this method incorporates all coordinates of the confusion matrix, MCC is considered one of the crucial evaluation metrics, particularly for imbalanced datasets [47].
@@ 4.5 Comparison with Existing Works
The majority of the research works have leveraged pretrained convolutional neural networks for classification. Among the well-known classifiers, EffecientNet and DenseNet are the most commonly used classifiers in classifying benign and malignant cancers [48, 2]. Through accurate fine-tuning and hyperparameter tuning, these models exhibit a noteworthy classification accuracy. Carcagni et al. [49] presented a classifier that integrates Support Vector Machine (SVM), a machine learning classifier, with DenseNet, a deep neural network. The modified network is more explainable and resource efficient. However, it failed to achieve a noticeable accuracy.
@@
EfficientNetV2-M [48] 0.9595 0.83 0.94 0.88
Custom deep neural network [50] 0.9016 0.9463 0.9391 0.9427
DenseNet with SVM [49] 0.8525 0.91 0.9175 0.9137
@@
Due to the large size of state-of-the-art image classifiers, some researchers have experimented with custom shallow neural networks which have produced impressive accuracy [50, 51, 52]. A comparative analysis of the proposed model with other research works on HAM10000 and Kaggle dataset is presented in Table 5 and 6 respectively. According to the analysis, the proposed solution achieves a superior performance over the majority of the research works.
@@
Modified DenseNet201 [2] 0.955 0.9702 0.9696 0.9546
Custom shallow CNN [51] 0.9874 0.9856 0.9927 0.997
SkinNet-16 [52] 0.9919 - - -

### v2

@@ 3.2.3 Highlighting ROI
Segmenting the lesion regions from the skin cancer images and leveraging the segmented portion for classification typically results in a better performance [bibi2022skin, monika2020skin]. However, incorporating a segmentation model like U-Net remarkably increases the computational burden. On the other hand, relying on traditional segmentation techniques might produce inconsistent results. Therefore, the lesion regions in the skin cancer images are highlighted to ensure feature preservation and prioritize relevant information. This step encompasses three key phases: grayscale conversion, image segmentation, and linear blending.
@@
Following the grayscale conversion, the image is then segmented with Otsu’s method. Otsu’s method calculates a threshold value based on the pixel intensity of the image which is leveraged for binary segmentation [peng2023multi]. For measuring the optimal threshold value, it iterates over the potential threshold values, and the value having the maximizing inter-class variance is returned as the threshold. The formula for inter-class variance is calculated using Equation 4. In this equation, $\omega_{0}(t)$ and $\omega_{1}(t)$ represent the probabilities of two classes separated by a threshold $t$. The equations of $\omega_{0}(t)$ and $\omega_{1}(t)$ are presented in Equation 5 and 6 respectively. $\omega_{0}(t)$ is the cumulative probabilities of pixel intensities up to the threshold $t$ while $\omega_{1}(t)$ measures the probabilities from $t+1$ to $L-1$, where $L-1$ is the highest intensity level. In Equation 4, $\sigma_{0}$ and $\sigma_{1}$ are the variances of two classes.
@@ 3.2.4 Unsharp Masking
Unsharp masking is an image sharpening method that employs low pass filters to sharpen details and improve overall visibility [song2022unsharp]. It improves skin cancer images by accentuating high-frequency components corresponding to edges and tiny details. The image sharpening algorithm has three major steps. In the first step, the input image is convoluted with a low pass filter. Let $I$ define the input image and $f$ be a low pass filter. Now the resulting blurred image, $I_{b}$ is obtained by operation illustrated in Equation 9. The selection of the low pass filter plays a crucial role in this step. For the proposed solution, the kernel is automatically determined from the standard deviation values of 2.0 in both the X and Y directions.
@@ 3.2.6 Resizing
Image resizing is also another frequently employed image preprocessing technique, particularly in the context of transfer learning. Typically state-of-the-art image classifiers are trained on ImageNet, that have a resolution of 469$\times$387 [deng2009imagenet]. However, they are typically resized to 255$\times$255 or 224$\times$244 [magdy2023performance]. Therefore, resizing the images to 224$\times$224 typically yields a better performance. Moreover, lowering the image size significantly reduces the computation. Since the images in the Kaggle dataset are already resized to the specified size, only the images from the HAM1000 dataset undergo this stage.
@@ 3.4 Teacher Model Creation
We have constructed a robust teacher model made of three feature extractors for guiding the lightweight student model. As presented in Figure 4, the proposed teacher model is composed of two streams. The first stream is composed of two CNN feature extractors: ResNet152V2 [he2016deep] and ConvNeXtBase [liu2022convnet], two state-of-the-art image classifiers. Since the feature extraction process of transformer architecture is different from CNNs, the second stream incorporates ViT Base [dosovitskiy2020image], a robust transformer network. The features from these streams are fused to construct a strong teacher model.
ResNet152V2, a variant of the ResNet architecture, is specifically engineered to mitigate the challenge of the vanishing gradient problem. This issue arises during the training of deep networks, where gradients diminish significantly during backpropagation, leading to slowed or stalled learning progress. Residual networks address this issue through the integration of residual blocks. A residual block creates a shortcut or residual connection by adding the block’s input to its output. The architectural enhancement allows the network to learn the residual mapping, which facilitates the model’s ability to detect gradients and identity mappings during training. Due to the fine gradient preservation, it is possible to create deep neural networks without encountering the vanishing gradient problem. ResNet152V2 extends the original ResNet architecture by incorporating batch normalization that speeds up the training process [hwooi2022deep]. This model consists of 152 layers that effectively extract useful features from the image, making the model a reliable classifier.
@@
The first stream of the proposed teacher model concatenates the features extracted from the pretrained ResNet152V2 and ConvNeXtBase. The combined features are then passed to batch normalization and Squeeze and Excitation (SE) block. SE block is an attention mechanism that prioritizes the channels that are more important for classification [hu2018squeeze]. The block consists of three operations: squeezing, exciting, and recalibrating. In the squeeze operation, illustrated in Equation 13, the spatial dimension of a channel is reduced with global average pooling. Let $X_{c}$ denote a feature map with $H$ height and $W$ width. Now, the summary of the channel, denoted with scaler $Z_{c}$, is obtained by taking the average of the channel $X_{c}$.
@@
The second stream, on the other hand, consists of a ViT Base and a batch normalization layer. The features extracted from the first stream and the second stream are combined to construct the final fusion model. The fusion model then follows three blocks of dropout and dense layer. The first two dropout layer randomly drops 50% neurons and the last layer drops 20% neurons. The dropout layers prevent overfitting. The first three dense layers, on the other hand, consist of 256, 128, and 64 neurons which are activated with Scaled Exponential Linear Unit (SELU). The SELU activation function, known for its simplicity and effectiveness in preventing overfitting, is utilized in the final layer, comprising two neurons employing the softmax activation function [verma2023revisiting]. The final layer comprises two neurons utilizing the softmax activation function. The model has over 236 million parameters out of which 3.39 million parameters are trainable. The size of the model is 975.85 megabytes.
@@ 3.5 Student Model Creation
The model is trained with knowledge distillation, a technique in which a simple model, known as the student, is trained to replicate the predictions of a more complex model, known as the teacher model [islam2023toward]. Figure 6 presents an overview of the knowledge distillation process. The approach involves sharing knowledge from the teacher to the student through softened probabilities, which are adapted using two key hyperparameters: temperature and alpha. The temperature parameter regulates the level of softening, which influences the smoothness of the probability distribution, whilst the alpha parameter sets the weighting of hard and soft targets throughout the training phase. This enables the student model to benefit from the rich knowledge stored in the teacher’s predictions, resulting in a more efficient and accurate learning process.
@@ 3.6 Ablation Study
In order to determine the optimal hyperparameters, an extensive ablation study has been conducted on the student model. The outcomes of the study on the HAM10000 and the Kaggle datasets are presented in Table 3 and 4 respectively. The results depict a similar impact of the seven hyperparameters. According to the analysis, the feature conversion mechanism is the most influential one among the seven parameters compared. While converting 2D future maps into 1D feature vectors, global max pooling produces the best performance in both the dataset while global average pooling has resulted in a poor performance. However, optimizer, temperature, alpha, and number of epochs are some of the least important hyperparameters. On the HAM10000 dataset, Adafactor has exhibited superior performance, outperforming RSMprop, which achieved the least accuracy. However, on the Kaggle dataset, RSMprop has emerged as the best optimizer among the four optimizers compared. Due to the conflicting outcome, Adafactor has been selected as the optimizer due to its faster convergence [harrison2022closer]. Another conflicting outcome is produced with the temperature hyperparameter. On the HAM10000 dataset, the best classification performance is produced with temperature 1, and temperature 5 results in the worst performance. On the other hand, the best performance is achieved with a temperature value of 5 on the Kaggle dataset. In the proposed model, the temperature is set to 1, which implies that the soft targets from the teacher model are passed without any scaling. On both datasets, an alpha value of 0.5, a learning rate of 0.002, and a batch size of 64 have produced the best performance. Notably, the learning rate has been one of the most influential hyperparameters in this experiment. In both datasets, a learning rate of 0.1 has created an overfitted model while reducing the learning rate has gradually improved the accuracy.
@@ 3.7 Model Quantization
Model quantization is a method that represents parameters and activations with a lower bit precision than the conventional 64-bit floating-point format, thereby reducing the memory footprint and computing demands of deep neural networks [rokh2023comprehensive]. This procedure is essential for deploying models on devices with limited memory and processing capability, like smartphones or edge devices. While quantization increases efficiency, over-quantization can reduce model accuracy due to the loss of fine-grained data [al2023unified].
@@ 4.2 Evaluation Metrics
To evaluate the model’s performance, we have considered five evaluation metrics including accuracy, precision, recall, F1-score and Matthews Correlation Coefficient (MCC). While accuracy presents the overall correctness of the model, precision measures the accuracy of positive predictions. Recall, on the other hand, the ratio of true positive predictions to the actual positives. Accuracy measures the harmonic mean of precision and recall, presenting a more balanced result. MCC considers true positive, true negative, false positive, and false negative values to asses the models’s performance. Since this method incorporates all coordinates of the confusion matrix, MCC is considered one of the crucial evaluation metrics, particularly for imbalanced datasets [hicks2022evaluation].
@@ 4.5 Comparison with Existing Works
The majority of the research works have leveraged pretrained convolutional neural networks for classification. Among the well-known classifiers, EffecientNet and DenseNet are the most commonly used classifiers in classifying benign and malignant cancers [venugopal2023deep, zia2022classification]. Through accurate fine-tuning and hyperparameter tuning, these models exhibit a noteworthy classification accuracy. Carcagni et al. [carcagni2019classification] presented a classifier that integrates Support Vector Machine (SVM), a machine learning classifier, with DenseNet, a deep neural network. The modified network is more explainable and resource efficient. However, it failed to achieve a noticeable accuracy.
@@
EfficientNetV2-M [venugopal2023deep] 0.9595 0.83 0.94 0.88
Custom deep neural network [ali2021enhanced] 0.9016 0.9463 0.9391 0.9427
DenseNet with SVM [carcagni2019classification] 0.8525 0.91 0.9175 0.9137
@@
Due to the large size of state-of-the-art image classifiers, some researchers have experimented with custom shallow neural networks which have produced impressive accuracy [ali2021enhanced, montaha2022shallow, ghosh2022skinnet]. A comparative analysis of the proposed model with other research works on HAM10000 and Kaggle dataset is presented in Table 5 and 6 respectively. According to the analysis, the proposed solution achieves a superior performance over the majority of the research works.
@@
Modified DenseNet201 [zia2022classification] 0.955 0.9702 0.9696 0.9546
Custom shallow CNN [montaha2022shallow] 0.9874 0.9856 0.9927 0.997
SkinNet-16 [ghosh2022skinnet] 0.9919 - - -